### v2

@@ 2.2.1 Darcy equation with cubic correction
where $b$ is a non-negative coefficient with the units $T^{2}/L^{2}$. Koch and Ladd (1997) simulated flow through periodic and random arrays of cylinders and confirmed that the first correction to Darcy’s law is cubic with respect to the bulk velocity. Firdaouss et al (1997) showed that the cubic correction also holds for anisotropic media provided that the modulus of the bulk velocity does not change under a reversal of the driving force. Hill et al (2001) investigated flow through random and ordered arrays of spheres using DNS and concluded that “[a]t all solid volume fractions, the first inertial contribution to the non-dimensional drag force was found to be proportional to the square of the Reynolds number, as predicted by the theory of Mei & Auriault.”.
@@ 2.2.4 Extended dynamic permeability model
The parameter $\xi\,\,[T/L]$ describes the nonlinearity and is related to the Forchheimer coefficient $b$ as $\xi=\frac{bK}{\mu}$. This model can be interpreted as an additive combination of the drag due to the boundary layers (corresponding to the high-frequency asymptotics (10) of Johnson et al (1987)) and of the drag in the steady state described by the Forchheimer equation (15).
@@ 2.3.3 Lower bounds for the coefficients of the unsteady Forchheimer equation
where the first term on the right hand side is the power added to or removed from the flow by the macroscopic pressure gradient. Comparing this equation to the volume-averaged kinetic energy equation (Zhu et al, 2014)
@@
wherein $\mathsf{S}$ is the strain rate tensor, we find that the term on the left hand side of (24) is the time derivative of a positive quantity and can take both signs. Thus it cannot be part of the dissipation. On the other hand, the second and third term on the right hand side of (24) have a definite sign. Hence they cannot be part of the time derivative of the kinetic energy which can take both signs. Consequently, we identify
@@ 3.2 Description of the simulation database
and therefore depends on two dimensionless parameters: The Hagen number $\mbox{{Hg}}=\left|g_{x}\right|d^{3}/(\rho\nu^{2})$ determines the amplitude and the Womersley number $\mbox{{Wo}}=\sqrt{\Omega\,d^{2}/\nu}$ determines the frequency of the macroscopic pressure gradient. The Reynolds number was defined as
@@
and was obtained as a result of the simulations. The simulations were performed at the three Womersley numbers $\mbox{{Wo}}=10$, $31.62$, and $100$ that correspond to the low, medium and high frequency regime in linear flow, respectively.222 The medium frequency regime in linear flow through a hexagonal sphere pack occurs around a Womersley number $\mbox{{Wo}}_{0}=\sqrt{\epsilon\,d^{2}/(\alpha_{\infty}\,K)}=30.5$, where $\Omega_{0}=\epsilon\,\nu/(\alpha_{\infty}\,K)$ is the transition frequency (Pride et al, 1993). The geometric constants are given in table 2. The Hagen numbers were set such that the simulations cover linear flow, laminar nonlinear flow, and transitional and turbulence-like flow at each Womersley number. The flow was started from rest and simulated until a recurrent behaviour could be observed in the superficial velocity. The difference in the superficial velocity in the last simulated cycle and the preceding cycle was less than approximately $0.1\%$ of the maximum superficial velocity in the last simulated cycle for the laminar cases and about $2\%$ for the transitional and turbulence-like cases. The simulation parameters can be found in table 1.
@@
The time series of the superficial velocity for some of the simulation cases is shown in figure 3. The amplitudes of the superficial velocity decrease with increasing Reynolds or Womersley number. The decrease of the amplitude with Womersley number is caused by the inertia of the system, whereas the decrease with Reynolds number can be explained by the increase of the drag. As the Reynolds number increases, the behaviour of the superficial velocity becomes non-sinusoidal. This could be explained by the effect of nonlinearity and turbulence. At high Womersley and Reynolds numbers, we can distinguish phases of strong and weak acceleration that lead to distinctive bends in the superficial velocity. In (Unglehrt and Manhart, 2023a) we could link this to an increase in the convective pressure drag due to flow separation at the contact points. As the Womersley number increases, we observe a phase lag of the superficial velocity with respect to the forcing. This can be seen from the zero-crossings of the superficial velocity which do not coincide with the zeros of the forcing ($\varphi=k\pi$ for $k=0,1,2,\dots$). This is due to the increasing importance of the inertia compared to the drag. On the other hand, the phase lag decreases as the Reynolds number increases. This goes in hand with an increase of the nonlinear drag which advances the inertial term. Overall, these time series demonstrate a distinct effect of nonlinearity and turbulence on the response of the superficial velocity to harmonic excitations.
@@ 3.3 Model constants for the hexagonal sphere pack
In this section, we specify the model constants that were used to evaluate the model predictions. The parameters for the linear models have rigorous definitions in terms of the steady Stokes flow ($K$, $\alpha_{0}$) and the potential flow solution ($\Lambda/d$, $\alpha_{\infty}$), respectively. Notably, these quantities are intrinsic properties of the pore geometry. The low frequency properties of the hexagonal sphere pack were determined from direct numerical simulations in the studies of Zhu and Manhart (2016) and Sakai and Manhart (2020); the high frequency properties were determined by a potential flow calculation in our previous study (Unglehrt and Manhart, 2023a). The parameter values are summarised in table 2.
@@
It can be seen in figure 4 that these coefficients provide a good fit to the simulation data. Finally, the nonlinear coefficient $\xi$ in the equations (18) and (23) results as $0.0154\,d/\nu$.
@@ 4.1 Comparison of model errors
In this section, we compare the accuracy of the model predictions in response to the forcing (30) with respect to the direct numerical simulation dataset. In particular, we look at the model of Turo and Umnova (2013) given in equation (18), the extended Pride et al (1993) model given in equation (23) and the unsteady Forchheimer equation (16). For the unsteady Forchheimer equations, we take the acceleration coefficient as the static viscous tortuosity $\alpha_{0}$ and as the high-frequency limit of the dynamic tortuosity $\alpha_{\infty}$.
Table 3: Amplitude and phase contribution to the $L^{2}$ error normalised with $\max\left\langle u\right\rangle_{\mathrm{s}}$ of the respective flow case. The entries where a linear model would be applied to a nonlinear flow case are marked as n.a. (not applicable).
@@
The error of the model predictions with respect to the direct numerical simulation results is quantified using the cycle-averaged $L^{2}$ error (37); we further decompose this error into an amplitude (41) and a phase contribution (42) using a Fourier series. Table 3 shows the amplitude and phase contribution to the $L^{2}$ error of the different model predictions with respect to the direct numerical simulation dataset. Note that the linear models were not applied to the nonlinear cases since they cannot account for the nonlinear drag and thus would produce large errors at high Reynolds numbers. Both the amplitude error and the phase error make significant contributions to the overall error of the models. The two components are weakly correlated, with the phase error being on average about $57\%$ of the amplitude error, but the importance of the two components varies considerably between models and simulation cases.
@@ 4.2 Analysis of prediction errors
For the turbulent case, the amplitude of the superficial velocity in the direct numerical simulation is $16\%$ lower than the amplitude of the linear case HF1 scaled by a factor of $10^{3}$ to the same Hagen number. Moreover, the maximum superficial velocity in the case HF6 is attained significantly earlier than in the case HF1 ($\varphi=0.61\pi$ for HF6 compared to $\varphi=0.94\pi$ for HF1). It can be seen in figure 7 that all models underpredict the amplitude of the superficial velocity: The error in the maximum superficial velocity is $-23\%$ for the model of Turo and Umnova (2013), $-21.2\%$ for the extended Pride et al (1993) model, $-26.4\%$ for the unsteady Forchheimer equation with $\alpha_{0}$ and $-12.9\%$ for the unsteady Forchheimer equation with $\alpha_{\infty}$. The phase is captured satisfactorily by the extended Pride et al (1993) model, the model of Turo and Umnova (2013) and the unsteady Forchheimer equation with $\alpha_{\infty}$, whereas the unsteady Forchheimer equation with $\alpha_{0}$ mispredicts the phase. It can be seen that all models fail to represent the relatively sharp bend at the peak.
@@
that is identical to the drag in the steady flow (15). For higher values of the acceleration coefficient, the drag becomes a double-valued function of the superficial velocity. Since the drag in the direct numerical simulation is also multi-valued, a nonzero virtual mass ($c>\rho/\epsilon$) is required to represent the correct behaviour. It can be seen that the peak superficial velocity of the model predictions decreases as the acceleration coefficient increases. The peak superficial velocity of the simulation cannot be reached even for $c=\rho/\epsilon$ for which highest amplitude of the different model predictions is obtained. As $b\left|\left\langle u\right\rangle_{\mathrm{s}}\right|\approx 12a$, this indicates that the nonlinear drag is too large.
@@ 5 Conclusion
In nonlinear flow, we compared the unsteady Forchheimer equation for two different choices of the acceleration coefficient, the extended dynamic permeability model of Turo and Umnova (2013) and an analogous formulation based on the model of Pride et al (1993). The unsteady Forchheimer equation with the acceleration coefficient based on the static viscous tortuosity $\alpha_{0}$ shows good results in the low frequency regime, but the results deteriorate at higher frequencies and are particularly bad in the high frequency nonlinear regime. On the other hand, the unsteady Forchheimer equation with the acceleration coefficient chosen based on the high-frequency limit of the dynamic tortuosity $\alpha_{\infty}$ shows very good results at high Reynolds numbers, but incurs large errors in the linear regime. The model of Turo and Umnova (2013) has relatively large errors throughout the entire parameter space. On the other hand, the extension of the model of Pride et al (1993) shows excellent results in the linear regime and at low frequencies numbers, but the results deteriorate for nonlinear flow at high frequencies.

### v3

@@ 2.2.1 Darcy equation with cubic correction
where $b$ is a non-negative coefficient with the units $T^{2}/L^{2}$. Koch and Ladd (1997) simulated flow through periodic and random arrays of cylinders and confirmed that the first correction to Darcy’s law is cubic with respect to the bulk velocity. Firdaouss et al (1997) showed that the cubic correction also holds for anisotropic media provided that the modulus of the bulk velocity does not change under a reversal of the driving force. Hill et al (2001) investigated flow through random and ordered arrays of spheres using DNS and concluded that “[a]t all solid volume fractions, the first inertial contribution to the non-dimensional drag force was found to be proportional to the square of the Reynolds number, as predicted by the theory of Mei & Auriault.”. Numerical investigations of Lasseux et al (2011) suggest that the cubic correction is valid for permeability Reynolds numbers $\mbox{{Re}}_{K}=\left|\left\langle\boldsymbol{u}\right\rangle_{\mathrm{s}}\right|\sqrt{K}/\nu$ up to about $0.1-0.8$ depending on the porosity and the pore geometry.
@@ 2.2.4 Extended dynamic permeability model
The parameter $\xi\,\,[T/L]$ describes the nonlinearity and is related to the Forchheimer coefficient $b$ as $\xi=bK/\mu$. This model can be interpreted as an additive combination of the drag due to the boundary layers (corresponding to the high-frequency asymptotics (10) of Johnson et al (1987)) and of the drag in the steady state described by the Forchheimer equation (15).
@@ 2.3.3 Lower bounds for the coefficients of the unsteady Forchheimer equation
where the first term on the right hand side is the power per unit volume added to or removed from the flow by the macroscopic pressure gradient. Comparing this equation to the volume-averaged kinetic energy equation (Zhu et al, 2014)
@@
wherein $\mathsf{S}$ is the strain rate tensor, we find that the term on the left hand side of (24) is the time derivative of a positive quantity. Hence, it can take both signs and may be arbitrarily large. Therefore, it cannot be part of the dissipation. On the other hand, if parts of the second and third terms on the right hand side of (24) belonged to the time derivative of the kinetic energy, then for a steady flow the kinetic energy would increase with time, which contradicts the assumption of a steady flow. Consequently, we identify
@@ 3.2 Description of the simulation database
and therefore depends on two dimensionless parameters: The Hagen number $\mbox{{Hg}}=\left|g_{x}\right|d^{3}/(\rho\nu^{2})$ determines the amplitude and the Womersley number $\mbox{{Wo}}=\sqrt{\Omega\,d^{2}/\nu}$ determines the frequency of the macroscopic pressure gradient. Geometrically, the Hagen number can be seen as the ratio of the distance over which the pressure gradient can move a particle within a viscous time unit $d^{2}/\nu$ to the sphere diameter $d$. Meanwhile, the Womersley number can be seen as the ratio of the sphere diameter to the thickness of the oscillatory boundary layer $\sqrt{\nu/\Omega}$. The Reynolds number was defined as
@@
and was obtained as a result of the simulations. The simulations were performed at the three Womersley numbers $\mbox{{Wo}}=10$, $31.62$, and $100$ that correspond to the low, medium and high frequency regime in linear flow, respectively.222 The medium frequency regime in linear flow through a hexagonal sphere pack occurs around a Womersley number $\mbox{{Wo}}_{0}=\sqrt{\Omega_{0}d^{2}/\nu}=\sqrt{\epsilon\,d^{2}/(\alpha_{\infty}\,K)}=30.5$, where $\Omega_{0}=\epsilon\,\nu/(\alpha_{\infty}\,K)$ is the transition frequency (Pride et al, 1993). The geometric constants are given in table 2. The Hagen numbers were set such that the simulations cover linear flow, laminar nonlinear flow, and transitional and turbulence-like flow at each Womersley number. The flow was started from rest and simulated until a recurrent behaviour could be observed in the superficial velocity. The difference in the superficial velocity in the last simulated cycle and the preceding cycle was less than approximately $0.1\%$ of the maximum superficial velocity in the last simulated cycle for the laminar cases and about $2\%$ for the transitional and turbulence-like cases. The simulation parameters can be found in table 1.
@@
The time series of the superficial velocity for some of the simulation cases is shown in figure 3. The amplitudes of the superficial velocity decrease with increasing Reynolds or Womersley number. The decrease of the amplitude with Womersley number is caused by the inertia of the system, whereas the decrease with Reynolds number can be explained by the increase of the drag. As the Reynolds number increases, the behaviour of the superficial velocity becomes non-sinusoidal. This could be explained mainly by the effect of nonlinearity (Unglehrt and Manhart, 2023a). Turbulence becomes significant only for the cases HF6 and HF7 (Unglehrt and Manhart, 2023b). At high Womersley and Reynolds numbers, we can distinguish phases of strong and weak acceleration that lead to distinctive bends in the superficial velocity. In (Unglehrt and Manhart, 2023a) we could link this to an increase in the convective pressure drag due to flow separation at the contact points. As the Womersley number increases, we observe a phase lag of the superficial velocity with respect to the forcing. This can be seen from the zero-crossings of the superficial velocity which do not coincide with the zeros of the forcing ($\varphi=k\pi$ for $k=0,1,2,\dots$). This is due to the increasing importance of the inertia compared to the drag. On the other hand, the phase lag decreases as the Reynolds number increases. This goes in hand with an increase of the nonlinear drag which advances the inertial term. Overall, these time series demonstrate a distinct effect of nonlinearity and turbulence on the response of the superficial velocity to harmonic excitations.
@@ 3.3 Model constants for the hexagonal sphere pack
In this section, we specify the model constants that were used to evaluate the model predictions. The parameters for the linear models have rigorous definitions in terms of the steady Stokes flow ($K$, $\alpha_{0}$) and the potential flow solution ($\Lambda/d$, $\alpha_{\infty}$), respectively. Notably, these quantities are intrinsic properties of the pore geometry. The low frequency properties of the hexagonal sphere pack were determined from direct numerical simulations in the studies of Zhu and Manhart (2016) and Sakai and Manhart (2020); the high frequency properties were determined by a potential flow calculation in our previous study (Unglehrt and Manhart, 2023a). The parameter values are summarised in table 2. The permeability value $1.731\cdot 10^{-4}\,d^{2}$ agrees well with the Kozeny-Carman correlation ($1.771\cdot 10^{-4}\,d^{2}$) and the self-consistent estimate of Boutin and Geindreau (2010) ($K_{c}=1.769\cdot 10^{-4}\,d^{2}$).
@@
It can be seen in figure 4 that these coefficients provide a good fit to the simulation data. The values also agree well with the Ergun equation for packed beds (Macdonald et al, 1979), which predicts $a=5647\,\mu/d^{2}$ and $b=76.3\,\rho/d$. Finally, the nonlinear coefficient $\xi$ in the equations (18) and (23) was calculated as $bK/\mu=0.0154\,d/\nu$.
@@ 3.4 Implementation of the models
In this section, we describe and verify the discretisation schemes that were used to evaluate the model predictions. The full derivation of these discretisations is given in (Unglehrt, 2024, appendix B.2).
@@ 4.1 Comparison of model errors
In this section, we compare the accuracy of the model predictions in response to the forcing (30) with respect to the direct numerical simulation dataset. In particular, we look at the model of Turo and Umnova (2013) given in equation (18), the extended Pride et al (1993) model given in equation (23) and the unsteady Forchheimer equation (16). For the unsteady Forchheimer equations, we take the acceleration coefficient as the static viscous tortuosity $\alpha_{0}$ and as the high-frequency limit of the dynamic tortuosity $\alpha_{\infty}$. For reference, we also include the unsteady Darcy equation (5) and the models of Johnson et al (1987) and Pride et al (1993) given by equation (13), where for the model of Johnson et al (1987) the relation (19) was used.
Table 3: Amplitude and phase contribution to the $L^{2}$ error normalised with $\max\left\langle u\right\rangle_{\mathrm{s}}$ of the respective flow case. The entries where a linear model would be applied to a nonlinear flow case are marked with a white background.
@@
The error of the model predictions with respect to the direct numerical simulation results is quantified using the cycle-averaged $L^{2}$ error (37); we further decompose this error into an amplitude (41) and a phase contribution (42) using a Fourier series. Table 3 shows the amplitude and phase contribution to the $L^{2}$ error of the different model predictions with respect to the direct numerical simulation dataset. Both the amplitude error and the phase error make significant contributions to the overall error of the models. The two components are weakly correlated, with the phase error being on average about $57\%$ of the amplitude error, but the importance of the two components varies considerably between models and simulation cases.
@@ 4.2 Analysis of prediction errors
For the turbulent case, the amplitude of the superficial velocity in the direct numerical simulation is $16\%$ lower than the amplitude of the linear case HF1 scaled by a factor of $10^{3}$ to the same Hagen number. Moreover, the maximum superficial velocity in the case HF6 is attained significantly earlier than in the case HF1 ($\varphi=0.61\pi$ for HF6 compared to $\varphi=0.94\pi$ for HF1). It can be seen in figure 7 that all models underpredict the amplitude of the superficial velocity: The error in the maximum superficial velocity is $-23\%$ for the model of Turo and Umnova (2013), $-21.2\%$ for the extended Pride et al (1993) model, $-26.4\%$ for the unsteady Forchheimer equation with $\alpha_{0}$ and $-12.9\%$ for the unsteady Forchheimer equation with $\alpha_{\infty}$. The phase is captured satisfactorily by the extended Pride et al (1993) model, the model of Turo and Umnova (2013) and the unsteady Forchheimer equation with $\alpha_{\infty}$, whereas the unsteady Forchheimer equation with $\alpha_{0}$ mispredicts the phase. This is reflected clearly in the phase error (table 3). Furthermore, figure 7 shows that all models fail to represent the relatively sharp bend at the peak.
@@
that is identical to the drag in the steady flow (15). For higher values of the acceleration coefficient, the drag becomes a double-valued function of the superficial velocity. Since the drag in the direct numerical simulation is also multi-valued, a nonzero virtual mass ($c>\rho/\epsilon$) is required to represent the correct behaviour. It can be seen that the peak superficial velocity of the model predictions decreases as the acceleration coefficient increases. The peak superficial velocity of the simulation cannot be reached even for $c=\rho/\epsilon$ for which the highest amplitude of the different model predictions is obtained. As $b\left|\left\langle u\right\rangle_{\mathrm{s}}\right|\approx 12a$, this indicates that the nonlinear drag is too large.
@@ 5 Conclusion
In nonlinear flow, we compared the unsteady Forchheimer equation for two different choices of the acceleration coefficient, the extended dynamic permeability model of Turo and Umnova (2013) and an analogous formulation based on the model of Pride et al (1993). The unsteady Forchheimer equation with the acceleration coefficient based on the static viscous tortuosity $\alpha_{0}$ shows good results in the low frequency regime, but the results deteriorate at higher frequencies and are particularly bad in the high frequency nonlinear regime. On the other hand, the unsteady Forchheimer equation with the acceleration coefficient chosen based on the high-frequency limit of the dynamic tortuosity $\alpha_{\infty}$ shows very good results at high Reynolds numbers, but incurs large errors in the linear regime. The model of Turo and Umnova (2013) has relatively large errors throughout the entire parameter space. On the other hand, the extension of the model of Pride et al (1993) shows excellent results in the linear regime and at low frequencies, but the results deteriorate for nonlinear flow at high frequencies.